# Few-Shot Learner Generalizes Across AI-Generated Image Detection

### v1

@@ Abstract
Current fake image detectors trained on large synthetic image datasets perform satisfactorily on limited studied generative models. However, they suffer a notable performance decline over unseen models. Besides, collecting adequate training data from online generative models is often expensive or infeasible. To overcome these issues, we propose Few-Shot Detector (FSD), a novel AI-generated image detector which learns a specialized metric space to effectively distinguish unseen fake images by utilizing very few samples. Experiments show FSD achieves state-of-the-art performance by $+7.4\%$ average ACC on GenImage dataset. More importantly, our method is better capable of capturing the intra-category common features in unseen images without further training.
@@ 1 Introduction
The development of generative models has led to significant strides in synthesizing photorealistic images, making it much difficult to distinguish AI-generated images from real ones. Over the past few years, diffusion-based models (Ho et al., 2020) have exceeded GANs (Goodfellow et al., 2014) and become the main stream paradigm of image generation, due to their strong ability of synthesizing images of high-quality. With the rapid release of open-source models and web APIs, users can easily create unrealistic images with brief texts. However, the increasing abuse of deepfake technology has dramatically influenced the human society and sparked unprecedented concerns on the credibility of online information.
@@
To curb the spread of AI-generated images used for malicious purposes, there is a growing urgency to develop detection methods for distinguishing fake images from authentic ones. Early works (Zhang et al., 2019b; Chai et al., 2020) focus on identifying artifacts in images, which prove to be particularly useful for detecting GAN-generated images. With the improvement of synthetic image datasets (Zhu et al., 2023; Boychev & Cholakov, 2024), detecting fake images from seen models is no longer difficult. Thus, a new challenge is to develop a classifier that generalizes across diverse unseen generative models. To address this problem, some studies (Ojha et al., 2023; Liu et al., 2024) adopt a large vision-language model to analyze images from a semantic perspective, while others (Wang et al., 2023; Luo et al., 2024) leverage the reconstruction ability of diffusion models.
@@
Another difficulty is the acquisition of training data. It is often expensive or infeasible to collect enough images from closed-source models such as DALL-E (Ramesh et al., 2021) and Midjourney (mid, 2022), due to their service prices and access restrictions. However, most existing fake image detectors (Wang et al., 2023; Ojha et al., 2023) require large amounts of training data to achieve good generalization, while fine-tuning without adequate data often results in overfitting. Furthermore, the ongoing release and updating of generative models pose a significant challenge for any trained detectors to keep up.
@@
After gaining a deep insight into these limitations, we propose a distinctive scheme based on few-shot learning to overcome them and design a comprehensive model to validate the efficacy of our solution. We first point out that current AI-generated image detection is a domain generalization task. Previous studies have been dedicated to finding a common indicator effective for all generative images. ̵However, they overlook the significant distinctions among data from different domains. We observe that images from unseen domains can actually be obtained in many real-world scenarios. Based on this fact, the complicated task can be transformed to a slightly easier one, called few-shot classification, by using relative few samples from unseen domains. Thus, we can extract rich domain information from these samples and use it to generalize across the unseen data.
@@
Based on the above analysis, we propose Few-Shot Detector (FSD), a novel method which can detect fake images with few given samples. In the few-shot classification task, synthetic images from different generative models are classified into separate categories, while the real images are classified into one class. The goal is to train a classifier capable of generalizing to unseen classes not in the training set, with only a small number of examples from them. To solve this problem, FSD utilizes a prototypical network (Snell et al., 2017), which learns a metric space to be able to handle unseen data. Different from end-to-end classification methods, FSD requires receiving a few samples from several classes to compute the corresponding prototypical representations. The test image is then classified with the nearest-neighbor method. Thus, the bias in unseen classes which can not be learned from the training set can now be automatically applied to the classifier without extra training.
@@
Experiments show that FSD obtains notable improvements over the state-of-the-art method by $+7.4\%$ average accuracy on GenImage dataset. We observe that only $10$ samples from the unseen classes can significantly improve the detection performance, which shows great privilege of our method. We also show that our approach is better equipped to keep pace with the rapid advancements in image generation, thereby highlighting the effectiveness of categorizing generated images by their sources. Our main contributions are as follows:
@@
  • To the best of our knowledge, we are the first to reconceptualize AI-generated image detection as a few-shot classification task, bringing it closer to the real-world scenario applications.
@@
  • We introduce an innovative synthetic image detection method, named Few-Shot Detector (FSD). Our approach is able to utilize few samples from unseen domains to achieve better detection performance.
@@ 2.1 Image Synthesis
Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) have revolutionized the field of generative modeling over the last decade. By simultaneously training a generator and a discriminator in an adversarial manner, GANs can generate samples of high resolution and good quality (Karras et al., 2020; Brock et al., 2019). However, they are prone to mode collapse, making it difficult for them to capture the full data distribution. An alternative technical route is likelihood-based modeling, which estimates the parameters of data distribution by maximizing the likelihood function. Within this framework, variational autoencoders (VAEs) (Kingma & Welling, 2014) ensure a smooth latent space which facilitates interpolation and exploration of new samples. Meanwhile, normalizing flow models (Papamakarios et al., 2021) allow for exact computation of the log-likelihood, which enables them to capture more complex data distributions. However, neither of them can match the high-quality image generation capabilities of GANs.
Diffusion Models (Ho et al., 2020; Song et al., 2021) have recently achieved remarkable performances in image synthesis. By reversing a gradual noising process, diffusion models can generate an image from random noise with several denoising steps. Ho & Salimans (2022) introduce a classifier-free guidance that enables the generation of images from text descriptions, eliminating the need for training a separate classifier. Another significant advancement is latent diffusion model (LDM) (Rombach et al., 2022), which applies diffusion process in latent space with a powerful autoencoder. This approach greatly enhances the visual fidelity of synthetic images while substantially reducing the required computational resources. Based on vast amounts of training data and substantial computing resources, large latent diffusion models (Podell et al., 2024; Esser et al., 2024) have become the most potent technique in image generation and the main source of synthetic images at present.
Figure 1: Detection pipeline of Few Shot Detector. FSD first maps the real and fake images from support set to the metric space by prototypical network and calculates the representation of each type. A test image is then mapped into the same space and classified according to the principle of nearest neighbor.
@@ 2.2 AI-generated Image Detection
Model-agnostic detection method. Learning based methods have been applied to detecting fake images from GANs (Zhang et al., 2019b; Qian et al., 2020; Chai et al., 2020). They treat synthetic image detection as a binary classification task, aiming to find the synthetic clues in generated images. Many studies (Zhang et al., 2019a; Gragnaniello et al., 2021) have found that classifiers do not generalize over unseen generative models, bringing generalization ability to the primary research objective. Meanwhile, Wang et al. (2020) reveal that detectors for CNN-generated images can be surprisingly generalizable due to the common systematic flaws shared among CNNs. Based on this work, Liu et al. (2023) analyze multi-view features to learn a robust representation, while Tan et al. (2024b) focus on high-frequency information to learn the source-agnostic features.
However, these common cues disappear when it comes to diffusion models, leading to the poor performance of earlier methods when applied to them. Therefore, more attention has been paid to semantic level. For instance, Ojha et al. (2023) prove that the informative image features from a pretrained CLIP model (Radford et al., 2021) can successfully be used for detecting fake images. Other studies (Liu et al., 2024; Tan et al., 2024a) deeply integrate textual information with images, thereby further enhancing the generalization capability. Moreover, patch-based detection method (Chen et al., 2024) has been proposed to reduce computational resources and accelerate detection speed. This method extracts features from a few patches instead of the whole image, enabling more efficient discrimination between fake and real images.
Diffusion-based detection method. The capability of open-source diffusion models to reconstruct images has opened a new technological route. DIRE (Wang et al., 2023) first uses the reconstruction loss to train a classifier, based on the observation that synthetic images can be reconstructed better than real ones. Following this discovery, SeDID (Ma et al., 2023) analyzes the loss errors between noised and denoised features at specific steps of processing, while LARE2 (Luo et al., 2024) focuses on the single-step reconstruction result in latent space. Both of them are able to improve the detection speed by not fully completing the entire reconstruction process. FakeInversion (Cazenavette et al., 2024) uses a BLIP (Li et al., 2022) to obtain a text description of the image for better reconstruction, which is capable of detecting images from unseen high-fidelity generators. AEROBLADE (Ricker et al., 2024) evaluates the autoencoder (AE) in different latent diffusion models and presents a training-free method for detecting fake images by AE reconstruction errors. These approaches achieve great success on open-source generative models. However, the differences between various diffusion models still can not be ignored, which makes it particularly challenging for them to detect synthetic images produced by numerous generative models, especially those that are closed-source.
@@ 3 Method
Figure 2: Prototypical network process. During training, the optimal center of each class is learned by minimizing the distances between query samples and their corresponding centers. During testing, The test image is only classified between real and fake.
In this section, we first describe the few-shot classification task of synthetic image detection. Then we elaborate on our novel representation, namely Few-Shot Detector (FSD), as illustrated in Fig. 1.
3.1 Few-Shot Synthetic Image Detection
Few-shot classification is a specialized machine learning task, in which the classifier is challenged to generalize to entirely new classes that are not present in the training set, using only a limited number of examples from these novel classes. Different from previous studies that categorize all the synthetic images produced by different generative models into a single fake class, we classify the images based on their generators. That is, we treat each group of them as a distinct class. In contrast, all the authentic images are classified into a single class. During detection, the set of additional labeled images used to support the classification task is known as the support set, while the set of those to be tested is referred to as the query set. The goal of this task is to determine which category each image in the query set belongs to, based on the support set.
In detail, we are given a support set comprising samples from $N$ new classes $\mathcal{S}=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\cdots\cup\mathcal{S}_{N}$, where $\mathcal{S}_{i}=\{(\mathbf{x}_{i1},y_{i1}),(\mathbf{x}_{i2},y_{i2}),\ldots,(\mathbf{x}_{iK},y_{iK})\}$ is the image subset of class $i$. $K$ is the number of samples in each class. $\mathbf{x}_{ij}\in\mathbb{R}^{D}$ is an image feature of $D$ dimension and $y_{ij}=i$ is the corresponding label. Each unlabeled image in the query set will be assigned to one of the $N$ classes. And this is the so-called $N$-way $K$-shot question.
@@ 3.2 Few-Shot Detector
Our FSD is based on the prototypical network (Snell et al., 2017), an effective approach for addressing the few-shot question above. FSD aims at learning a metric space in which samples from the same class are close in distance, while samples from different classes are relatively distant from each other. Thus, we can represent each class by one single vector, known as the prototypical representation. For each sample in the query set, classification is then conducted by simply finding the nearest representation to it.
In detail, as depicted in Fig. 1, FSD first adopts a neural network $f_{\phi}:\mathbb{R}^{D}\to\mathbb{R}^{M}$ to compute the $M$-dimensional vector in the metric space of an image, where $\phi$ is the learnable parameters. A prototypical representation $\mathbf{c}_{i}$ is defined as the mean of embedded vectors from its support set $\mathcal{S}_{i}$:
$\mathbf{c}_{i}=\frac{1}{|\mathcal{S}_{i}|}\sum_{\mathbf{x}_{j}\in\mathcal{S}_{i}}f_{\phi}(\mathbf{x}_{j})$ (1)
After calculating all the $N$ prototypical representations, for an image $\mathbf{x}$ in the query set, we compute its vector $f_{\phi}(\mathbf{x})$ and the distances between it and all the representations $d(f_{\phi}(\mathbf{x}),\mathbf{c}_{i})$, where $d:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}_{0}^{+}$ is a distance function. The probability that the query sample $\mathbf{x}$ belongs to class $i$ is calculated by:
$p(y=i|\mathbf{x})=\mathrm{softmax}_{1\leq i\leq N}(-d(f_{\phi}(\mathbf{x}),\mathbf{c}_{i}))$ (2)
The image $\mathbf{x}$ is then assigned to the class with maximum probability, or the nearest prototypical representation to it.
During training, we randomly selected $N_{c}$ classes from the training set $\mathcal{D}_{train}$. For each selected class $i$, we randomly choose $N_{s}$ samples as support set $\mathcal{S}_{i}$ and another $N_{q}$ samples as query set $\mathcal{Q}_{i}$. We compute the $N_{c}$ prototype centers $\mathbf{c}_{i}$ according to equation 1. The prototypical network outputs a probability distribution over the $N_{c}$ classes for each query sample $\mathbf{x}$ in query set. The optimization target is to minimize the the negative log-probability $J(\phi)$. Pseudo code to compute the loss $J(\phi)$ for a training step is provided in Algorithm 1.
@@
  Input: Training set $\mathcal{D}=\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\cdots\mathcal{D}_{N}$, where $\mathcal{D}_{k}$ denotes the subset of class $k$, number of training class $N_{C}$, number of support examples per class $N_{S}$, number of query examples per class $N_{Q}$.
  Output: training loss $J$ of one step.
  $V\leftarrow\textit{randomly select }N_{c}\textit{ classes from }\{1,2,\ldots,N\}$
@@
     $S_{k}\leftarrow\textit{randomly select }N_{s}\textit{ samples from }\mathcal{D}_{k}$
     $Q_{k}\leftarrow\textit{randomly select }N_{q}\textit{ samples from }\mathcal{D}_{k}/S_{k}$
@@
     $J\leftarrow J+\frac{1}{N_{c}N_{q}}\sum_{\mathbf{x}_{j}\in Q_{k}}\log\mathrm{softmax}(-d(f_{\phi}(\mathbf{x}_{j}),\mathbf{c}_{k}))$

### v2

@@ Abstract
Current fake image detectors trained on large synthetic image datasets perform satisfactorily on limited studied generative models. However, these detectors suffer a notable performance decline over unseen models. Besides, collecting adequate training data from online generative models is often expensive or infeasible. To overcome these issues, we propose Few-Shot Detector (FSD), a novel AI-generated image detector which learns a specialized metric space for effectively distinguishing unseen fake images using very few samples. Experiments show that FSD achieves state-of-the-art performance by $+11.6\%$ average accuracy on the GenImage dataset with only $10$ additional samples. More importantly, our method is better capable of capturing the intra-category commonality in unseen images without further training. Our code is available at https://github.com/teheperinko541/Few-Shot-AIGI-Detector.
@@ 1 Introduction
The development of generative models has led to significant strides in synthesizing photorealistic images, making it much more difficult to distinguish AI-generated images from real ones. Over the past few years, diffusion-based models (Ho et al., 2020) have exceeded GANs (Goodfellow et al., 2014) and become the mainstream paradigm of image generation, due to their exceptional ability to synthesize high-quality images. With the rapid release of open-source models and web APIs, users can easily create vivid images from brief textual descriptions. However, the increasing abuse of deepfake technology has dramatically influenced human society and sparked unprecedented concerns on the credibility of online information.
@@
To curb the spread of AI-generated images used for malicious purposes, there is a growing urgency to develop detection methods capable of distinguishing fake images from authentic ones. Early works (Zhang et al., 2019b; Chai et al., 2020) focus on identifying artifacts in synthetic images, which have proved to be particularly useful for detecting GAN-generated images. With the improvement of synthetic image datasets (Zhu et al., 2023; Boychev & Cholakov, 2024), detecting fake images from seen generative models is no longer a difficult task. Thus, a novel challenge arises in developing classifiers that generalize across unseen models. To address this problem, some studies (Ojha et al., 2023; Liu et al., 2024) adopt a large vision-language model to analyze images from a semantic perspective, while others (Wang et al., 2023; Luo et al., 2024) leverage the reconstruction ability of diffusion models.
@@
Another difficulty is the acquisition of training data. It is often expensive or infeasible to collect enough images from closed-source models such as DALL-E (Ramesh et al., 2021) and Midjourney (mid, 2022). This is due to their service prices and access restrictions. However, most existing fake image detectors (Wang et al., 2023; Ojha et al., 2023) require large amounts of training data to achieve good generalization, and fine-tuning without adequate data often results in overfitting. Furthermore, the ongoing release and updating of generative models pose a significant challenge for existing detection systems to keep up.
@@
After gaining a deep insight into these limitations, we propose an innovative approach based on few-shot learning to overcome them. A robust model is also developed to validate the efficacy of our solution. We first point out that current AI-generated image detection is a domain generalization task. Previous studies have been dedicated to discovering a universal indicator effective for detecting diverse fake images. ̵However, they overlook the significant distinctions among data from different domains. We observe that images from unseen domains can actually be obtained in many real-world scenarios. Based on this fact, the complicated task can be transformed into a more tractable one, called few-shot classification, through limited samples from unseen domains. Thus, we can extract rich domain-specific features from the given samples and use them to generalize across the unseen data, as shown in Figure 1.
Figure 1: Challenge facing previous works and our solution. Most classifiers exhibit significant performance degradation over unseen data. We address this challenge by introducing a few-shot strategy which is able to make full use of the limited unseen samples.
@@
Based on the above analysis, we propose Few-Shot Detector (FSD), a novel method which is able to detect fake images with few given samples. In the few-shot classification task, synthetic images from different generative models are classified into separate categories, whereas the real images are classified into a single category. The objective is to develop a classifier capable of generalizing to unseen classes that are not in the training set, with only a limited number of examples from them. To solve this problem, FSD utilizes the Prototypical Network (Snell et al., 2017), which learns a metric space to effectively handle unseen data. The given samples are used to compute the prototypical representations of different classes. The test image is then classified using the nearest-neighbor method by comparing it against these representations. Thus, FSD can automatically adjust for bias in previously unseen classes without additional training or fine-tuning.
@@
Experiments show that FSD achieves notable improvements over the state-of-the-art method by $+11.6\%$ average accuracy on the GenImage dataset. We observe that only $10$ samples from the tested classes can significantly improve the detection performance, which highlights the superiority and practicality of FSD. We also show that our approach is better equipped to keep pace with the rapid advancements in image generation, thereby demonstrating the effectiveness of categorizing generated images by their sources.
Our contributions are summarized as follows:
@@
  • To the best of our knowledge, we are the first to reconceptualize AI-generated image detection as a few-shot classification task, bringing this task closer to real-world applications.
@@
  • We introduce an innovative synthetic image detection method, named Few-Shot Detector (FSD). Our approach is able to utilize a few samples from unseen domains to achieve better detection performance.
@@ 2.1 Image Synthesis
Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) have revolutionized the field of generative modeling over the last decade. By simultaneously training a generator and a discriminator in an adversarial manner, GANs can generate samples of high resolution and good quality (Brock et al., 2019; Karras et al., 2020). However, they are prone to mode collapse, which makes it difficult for them to fully capture the data distribution. An alternative approach is likelihood-based modeling, which estimates the parameters of data distributions by maximizing the likelihood function. Based on this methodology, variational autoencoders (VAEs) (Kingma & Welling, 2014) ensure a smooth latent space that facilitates interpolation and exploration of new samples. In contrast, normalizing flow models (Papamakarios et al., 2021) enable exact log-likelihood computation, allowing them to capture more complex data distributions. However, neither of them can match GANs in high-quality image generation.
Diffusion Models (Ho et al., 2020; Song et al., 2021) have recently achieved remarkable performance in image synthesis. By reversing a gradual noising process, diffusion models can generate images from random noise through successive denoising steps. Ho & Salimans (2022) introduce classifier-free guidance which enables generating images from text descriptions, eliminating the need for training a separate classifier. Another significant advancement is the Latent Diffusion Model (LDM) (Rombach et al., 2022), which applies the diffusion process in latent space with a powerful autoencoder. This approach greatly enhances the visual fidelity of synthetic images while substantially reducing the required computational resources. Leveraging vast amounts of training data, large-scale latent diffusion models (Podell et al., 2024; Esser et al., 2024) have emerged as the leading technique in image generation and the main source of synthetic images at present.
Figure 2: Detection pipeline of Few-Shot Detector. FSD first maps the real and fake images from the support set to the metric space by the Prototypical Network and calculates the representation of each type. A test image is then mapped into the same space and classified according to the nearest-neighbor principle.
@@ 2.2 AI-generated Image Detection
Model-agnostic detection method. Learning-based methods have been applied to detecting fake images from GANs (Zhang et al., 2019b; Qian et al., 2020; Chai et al., 2020). They treat synthetic image detection as a binary classification task, aiming to find synthetic clues in generated images. Numerous studies (Zhang et al., 2019a; Gragnaniello et al., 2021) have found that classifiers often fail to generalize across unseen generative models, bringing generalization ability to the primary research objective. Wang et al. (2020) reveal that detectors for CNN-generated images can be surprisingly generalizable due to the common systematic flaws shared among CNNs. Based on this work, Liu et al. (2023) analyze multi-view features to learn a robust representation, while Tan et al. (2024b) focus on high-frequency information to capture source-agnostic features.
However, these common cues disappear when it comes to diffusion models, leading to the poor performance of earlier methods when applied to them. Therefore, more attention has been paid to semantic level. For instance, Ojha et al. (2023) prove that the informative image features from a pretrained CLIP model (Radford et al., 2021) can be successfully used for detecting fake images. Other studies (Liu et al., 2024; Tan et al., 2024a) deeply integrate textual information with images, thereby further enhancing the generalization capability. Moreover, a patch-based detection method (Chen et al., 2024) extracts features from a single patch rather than the entire image, which significantly accelerates the detection speed.
Diffusion-based detection method. The reconstruction capability of open-source diffusion models has opened a new technological pathway. DIRE (Wang et al., 2023) first uses the reconstruction loss to train a classifier, based on the observation that synthetic images can be reconstructed with higher fidelity compared to real ones. SeDID (Ma et al., 2023) analyzes the loss errors between noised and denoised features at specific steps of processing, while LARE2 (Luo et al., 2024) focuses on the single-step reconstruction result in latent space. Both of them improve the detection speed by not fully completing the entire reconstruction process. FakeInversion (Cazenavette et al., 2024) employs BLIP (Li et al., 2022) to obtain a text description of the image for better reconstruction, which is capable of detecting unseen high-fidelity generated images. AEROBLADE (Ricker et al., 2024) evaluates autoencoder (AE) in different latent diffusion models and presents a training-free method for detecting fake images based on AE reconstruction errors. Although these approaches have achieved great success on open-source generative models, they still face challenges in detecting synthetic images from various inaccessible models due to the substantial differences among them.
@@ 3 Method
In this section, we first describe the few-shot classification task for synthetic image detection. Then we elaborate on our novel representation, namely Few-Shot Detector (FSD), as illustrated in Figure 2.
3.1 Few-shot Synthetic Image Detection
Few-shot classification is a specialized machine learning task, where the classifier is challenged to generalize to entirely new classes not present in the training set, using only a limited number of examples from these novel classes. Different from end-to-end detection methods that aggregate all synthetic images produced by different generative models into a single fake class, we classify these images based on their generators. This means that we consider the productions of each model as an individual class. Conversely, authentic images continue to be grouped into a single class. During detection, the support set comprises labeled images aiding in the classification task, while the query set contains unlabeled images to be classified. The objective of this task is to determine the category of each image in the query set based on the support set.
In detail, we are given a support set $\mathcal{S}=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\cdots\cup\mathcal{S}_{N}$, which comprises samples from $N$ new classes. Each subset $\mathcal{S}_{i}=\{(\mathbf{x}_{i1},y_{i1}),(\mathbf{x}_{i2},y_{i2}),\ldots,(\mathbf{x}_{iK},y_{iK})\}$ denotes the collection of images from class $i$. $K$ is the number of samples in each class. $\mathbf{x}_{ij}\in\mathbb{R}^{D}$ is an image of $D$ dimensions and $y_{ij}=i$ indicates its corresponding label. Each unlabeled image in the query set will be assigned to one of the $N$ classes. This is the so-called $N$-way $K$-shot task.
@@ 3.1 Few-shot Synthetic Image Detection
Figure 3: Prototypical calculation. During training, the metric space is learned by minimizing the distances between query samples and their corresponding centers. During testing, the test image is only classified between real and fake.
@@ 3.2 Few-Shot Detector
Our FSD is based on the Prototypical Network (Snell et al., 2017), an effective method for addressing the few-shot question above. The Prototypical Network aims to learn a metric space in which samples from the same class are close in distance, while samples from different classes are relatively distant from each other. Thus, we can represent each class by a single vector, known as the prototypical representation. For each sample in the query set, classification is conducted by simply finding the nearest representation to it.
In detail, as depicted in Figure 2, FSD adopts a neural network $f_{\phi}:\mathbb{R}^{D}\to\mathbb{R}^{M}$ to compute the $M$-dimensional vector in the metric space of an image, where $\phi$ is the learnable parameters. The prototypical representation $\mathbf{c}_{i}$ for class $i$ is defined as the average of the embedded vectors derived from its corresponding support set $\mathcal{S}_{i}$:
$\mathbf{c}_{i}=\frac{1}{|\mathcal{S}_{i}|}\sum_{\mathbf{x}_{j}\in\mathcal{S}_{i}}f_{\phi}(\mathbf{x}_{j})\text{.}$ (1)
After gaining all the $N$ prototypical representations and the vector $f_{\phi}(\mathbf{x}_{q})$ of a test image $\mathbf{x}_{q}$ in the query set, we compute the distances between the test vector and all the representations $d(f_{\phi}(\mathbf{x}_{q}),\mathbf{c}_{i})$, using a distance function $d:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}_{0}^{+}$. The probability that the query sample $\mathbf{x}_{q}$ belongs to class $i$ is determined by:
$p(y=i|\mathbf{x}_{q})=\mathrm{Softmax}_{1\leq i\leq N}(-d(f_{\phi}(\mathbf{x}_{q}),\mathbf{c}_{i}))\text{.}$ (2)
The image $\mathbf{x}_{q}$ is assigned to the class with the maximum probability, or the nearest prototypical representation to it.
As shown in Figure 3(a), during each training step, $N_{c}$ classes are randomly selected from the training set. For each selected class $i$, we randomly choose $N_{s}$ samples from it as the support set $\mathcal{S}_{i}$ and another $N_{q}$ samples as the query set $\mathcal{Q}_{i}$. We compute the $N_{c}$ prototypical representations according to Equation 1. The Prototypical Network outputs a probability distribution over the $N_{c}$ classes for each sample in the query set. The optimization target is to minimize the negative log-probability $J(\phi)$. Pseudo code to compute the loss $J(\phi)$ is provided in Algorithm 1.
@@
  Input: Training set $\mathcal{D}=\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\cdots\mathcal{D}_{N}$, where $\mathcal{D}_{k}$ denotes the subset of class $k$; number of training class $N_{c}$; number of support examples per class $N_{s}$; number of query examples per class $N_{q}$.
  Output: training loss $J$.
  $V\leftarrow\text{randomly select }N_{c}\text{ classes from }\{1,2,\ldots,N\}$
@@
     $S_{k}\leftarrow\text{randomly select }N_{s}\text{ samples from }\mathcal{D}_{k}$
     $Q_{k}\leftarrow\text{randomly select }N_{q}\text{ samples from }\mathcal{D}_{k}/S_{k}$
@@
     $J\leftarrow J-\frac{1}{N_{c}N_{q}}\sum_{\mathbf{x}_{j}\in Q_{k}}\log\mathrm{Softmax}(-d(f_{\phi}(\mathbf{x}_{j}),\mathbf{c}_{k}))$